# Practical insights on the effect of different encodings, ansätze and measurements in quantum and hybrid convolutional neural networks

### v1

@@ Abstract
This study investigates the design choices of parameterized quantum circuits (PQCs) within quantum and hybrid convolutional neural network (HQNN and QCNN) architectures, applied to the task of satellite image classification using the EuroSAT dataset. We systematically evaluate the performance implications of data encoding techniques, variational ansätze, and measurement in $\approx 500$ distinct model configurations. Our analysis reveals a clear hierarchy of influence on model performance. For hybrid architectures, which were benchmarked against their direct classical equivalents (e.g. the same architecture with the PQCs removed), the data encoding strategy is the dominant factor, with validation accuracy varying over $30\%$ for distinct embeddings. In contrast, the selection of variational ansätze and measurement basis had a comparatively marginal effect, with validation accuracy variations remaining below $5\%$. For purely quantum models, restricted to amplitude encoding, performance was most dependent on the measurement protocol and the data-to-amplitude mapping. The measurement strategy varied the validation accuracy by up to $30\%$ and the encoding mapping by around $8$ percentage points.
@@
In recent years, quantum machine learning (QML) has emerged as an interdisciplinary field at the intersection of quantum computing and machine learning Schuld et al. (2014a); Wittek (2014); Biamonte et al. (2017); Schuld and Petruccione (2018). Its objective is to leverage phenomena in quantum mechanics to tackle problems from machine learning. A central component of many contemporary QML models is the parameterized quantum circuit (PQC), a flexible framework for creating trainable quantum circuits that can be used independently or integrated as hybrid components within classical neural networks (NNs).
This work investigates the use PQCs for the task of satellite image classification using the EuroSAT dataset Helber et al. (2019). These PQCs serve as trainable quantum components that can be integrated into NNs or used by themselves. While PQCs offer a flexible paradigm for hybrid and purely QML models, their efficacy is highly sensitive to a multitude of design choices, with the most important being: the encoding technique, the ansätze and the measurement strategy.
We study two different hybrid architectures networks and a purely quantum one, all inspired by convolutional neural networks (CNN). For the hybrid ones we have a “quantum first” model, in which PQCs are used as convolutions to extract features, and a “quantum later” model Senokosov et al. (2024), in which PQCs are used between the fully connected layers to process the extracted features Senokosov et al. (2024); Hur et al. (2022); Elhag et al. (2024). For the quantum architecture, we use three variations of PQCs as standalone components, one of which inspired by the QCNN, introduced in Cong et al. (2019).
Our approach builds on the work of Bowles et al. (2024), which systematically benchmarked different quantum machine learning models by varying the encoding, measurement and hyperparameters. Our difference lies in the fact that we fix the hyperparameters and benchmark our hybrid and purely quantum models across all three key circuit components: encoding, ansätze and measurement.
@@
All our experiments were done through simulations using the CTIC Quantum Testbed (QUTE) platform García Mangas and Alonso (2025). We used Pennylane Bergholm et al. (2022) to implement the PQCs, both in the purely quantum and hybrid models, and PyTorch P. et al. (2019) for the classical layers.
.1 Image classification
Our focus will be in classifying satellite images from the EuroSAT dataset Helber et al. (2019), examples can be seen in FIG. 1. The dataset consists of $27,000$ geo-referenced images, provided by the Copernicus Earth observation program European Space Agency (nd) using the Sentinel-2 satellite, and comes in two versions: RGB and all bands. The first consists of only RGB channels and the other uses all the bands of the spectrum as collected from Sentinel-2. Each image channel has a resolution of $64\times 64$ pixels.
In our experiments, we used the RGB version and downscaled them to $32\times 32$ and $16\times 16$. The former was only studied for the hybrid models, due computation constraints, and the latter for both hybrid and purely quantum models.
We selected this dataset as a continuation of the work conducted within the ARQA Cervera Technology Transfer Network Red ARQA (2024). Nevertheless, identical experiments can be replicated on alternative datasets, provided they satisfy the previously established size constraints.
Figure 1: Samples from the EuroSAT dataset introduced in Helber et al. (2019). Each sample is a $64\times 64\times 3$ RGB image obtained from the Sentinel-2 satellite.
.2 Parameterized quantum circuits
PQCs are the cornerstone of our approach. We integrate them as layers consisting of one or more circuits in the hybrid models and use them by themselves in the purely quantum ones. In general, PQCs consist of three key components: encoding, used to embed the information in a quantum state, variational ansatz, used to manipulate the encoded information, and measurement, used to extract information from the quantum state. FIG. 2 shows the prototypical example of a PQC.
[…]
@@
.2.1 Data encoding
As previously mentioned, the first stage in a PQC is the data encoding, this step serves as a mapping from data to a quantum state. Given that there is not a one size fits all solution for the encoding step, we chose to include the following ones into our analysis:
  • Angle encoding Schuld and Petruccione (2018); Schuld et al. (2014b); Cao et al. (2017): This is a prevalent and relatively straightforward method where $N$ classical data features, $x_{1},x_{2},\ldots,x_{N}$, are encoded into the rotation angles of $N$ individual qubits. Each feature $x_{j}$ is used as an angle parameter for a single-qubit rotation gate, such as $R_{X}(\theta_{j})$, $R_{Y}(\theta_{j})$, or $R_{Z}(\theta_{j})$, applied to a specific qubit. Notably, using $R_{Z}(\theta_{j})$ as an encoding gate acting on an initial initial state from the computational basis only adds an irrelevant global phase.
  • Amplitude encoding Schuld and Petruccione (2018): This technique embeds a classical data vector of $2^{n}$ normalized real numbers, $\{\alpha_{i}\}$, directly into the amplitudes of an $n$-qubit quantum state vector $|\Psi\rangle$. The state is represented as:
@@
    where $|i\rangle$ are the computational basis states, and the amplitudes must satisfy the normalization condition $\sum_{i=0}^{2^{n}-1}|\alpha_{i}|^{2}=1$. The state preparation process required to achieve an arbitrary state $|\Psi\rangle$ can be computationally expensive, potentially demanding an exponentially deep circuit Knill (1995).
@@
  • Instantaneous quantum polynomial-time (IQP) encoding Havlíček et al. (2019): This encoding extends the concept of angle encoding by incorporating entangling operations. $N$ classical features are typically encoded into $N$ qubits using single-qubit rotations, akin to angle encoding. These rotations are then interspersed with diagonal two-qubit gates, commonly $R_{ZZ}(\phi_{jk})$ gates, where $\phi_{jk}$ can be a function of individual features or products of features (e.g., $x_{j}x_{k}$). This structure allows for the generation of more complex quantum states that can potentially capture correlations between input features. The number of layers is a tunable parameter.
@@
  • Quantum approximate optimization algorithm (QAOA) inspired encoding Lloyd et al. (2020): This encoding strategy draws inspiration from the structure of the QAOA. It transforms $N$ classical features into parameters that define a QAOA-like ansatz applied to $N$ qubits. The features typically determine the angles for rotation gates within layers corresponding to a problem Hamiltonian and a mixer Hamiltonian. The number of layers is a tunable parameter and there are 3 variants: $X$, $Y$ and $Z$.
@@
  • Ring encoding Venturelli (2023): This method combines angle encoding for $N$ features on $N$ qubits with a specific entangling topology. Following initial single-qubit rotations that embed the classical data, two-qubit entangling gates are applied between adjacent qubits arranged in a ring structure. Specifically, qubit $i$ is entangled with qubit $(i+1)\mod N$. The number of layers is a tunable parameter.
@@
  • Waterfall encoding Venturelli (2023): Similar to Ring Encoding, Waterfall Encoding also utilizes angle encoding for $N$ features on $N$ qubits, followed by a structured application of entangling gates. In this scheme, the entangling gates are typically arranged in a “waterfall” pattern. For example, qubit $i$ might control an operation on qubit $i+1$, which in turn might control an operation on qubit $i+2$, and so forth.
.2.2 Variational ansatz
@@
The variational circuit constitutes the second part of a PQC. It consists of a set of quantum gates with trainable parameters, akin to neurons in a NN. Its purpose is to manipulate the quantum state generated by the previous step with the goal of turning it into a state whose measurement, in a given basis, solves the task in hand. We chose to experiment with the following ones:
@@
  • No entangled: This circuit uses randomness in the assignment of quantum gates, applying either an RZ gate or the identity operation to each qubit in a stochastic manner. Due to the potential presence of trivial operations (identity), the resulting circuit typically exhibits a lower computational depth compared to more densely parameterized or entangled circuits.
@@
  • Fully entangled Schuld et al. (2020): This circuit is implemented using the StronglyEntanglingLayers module from PennyLane Bergholm et al. (2022). It allows for a configurable number of layers, each consisting of a sequence of single-qubit rotations followed by entangling operations. This structure enables flexible depth and entanglement patterns suitable for a variety of variational quantum algorithms.
@@
  • Ring circuit Venturelli (2023): It performs single-qubit rotations using RY and RZ gates, followed by the application of controlled-Z (CZ) gates arranged in a ring topology. The resulting structure resembles the ring embedding, but with entanglement introduced through CZ operations instead of CNOT.
@@
  • NQ circuit Hur et al. (2024): The circuit consists of multiple layers, each combining parameterized single-qubit rotations, RY and RZ, followed by entangling gates (CNOT) in a linear nearest-neighbor topology.
@@
  • QCNN circuit Cong et al. (2019): circuit structure inspired by classical CNNs in which convolutional and pooling operations are subsequently applied in each layer. In our implementation, when the number of qubits at a layer is odd, one is excluded from the pooling, although it is still considered in the convolutional stages. At the end of the circuit, the qubits that were discarded are recovered for the fully connected layer. Given that the number of free paremeters in an arbitrary unitary scales as $14^{N}-$ 1, we substituted this last operation by the Simplified Two Design Cerezo et al. (2021).
.2.3 Measurement
@@
  • Z-basis: It is also called the computational basis, this is the most commonly used measurement in quantum circuits because it collapses a qubit into $|0\rangle$ or $|1\rangle$.
@@
  • Histogram: We measure each qubit on the Z basis and count the frequency of each of the possible $2^{N}$ outcomes, for $N$ qubits. Each state is associated with a class, hence the model in this case will try to learn to link classes with corresponding states. This is in essence a Quantum Circuit Born Machine Liu and Wang (2018) in which we set the most frequent sample to be the predicted class label.
  • Pauli string (Paulis): We estimate the expectation value of pauli strings of $N$ qubits, i.e., observables of the form: $O=\otimes_{i=0}^{N}P_{i}$, where $P_{i}$ is a Pauli gate acting on the $i-th$ qubit. Here, a class is identified with an operator such that the model is trained to maximize the expectation value of that observable when images from the correct class are fed into the model.
Recent studies have also pointed out the relevance of the type of measurement considered in the models Chen et al. (2025).
I Models
For an overview of the models see TABLE 1.
I.1 Hybrid models I: HQNN-Parallel
The HQNN-Parallel was firstly introduced in Senokosov et al. (2024) (shown in FIG. 4(a)). This architecture introduces NN layer based on multiple PQCs. From this point onward, we will call this a quantum linear layer due to its similarity with a classical linear layer.
A quantum linear layer operates in a similar manner as its classical counterpart. It takes in an input vector of size $V$ and outputs another of size $W$. However, unlike the classical case, here $W\leq V$ given that at most we do a single measurement on all qubits. In the classical version, the linear layer is defined by the number of input and output features. The quantum analog is characterized by the number of input features, the number of qubits allocated per circuit, same number for all, and the encoding used, which will determine the proportion between input and output features.
In the simplest case, we have an input vector of length $L$, we allocate $L/2$ qubits per circuit and use an encoding that maps a single feature to a single qubit, like angle encoding. In this case, the quantum linear layer will consists of two PQCs of $L/2$ qubits each, and the output will be a vector of length $L$, given that we measure all qubits on a given basis. This scheme is shown in FIG. 3. For instance, if we use amplitude embedding, then each $N$ qubit circuit will process $2^{N}$ features and the output will be of size $N\times\text{the number of circuits}$.

### v2

@@ Abstract
This study investigates the design choices of parameterized quantum circuits (PQCs) within quantum and hybrid convolutional neural network (HQNN and QCNN) architectures applied to the task of satellite image classification using the EuroSAT dataset. We systematically evaluate the performance implications of data encoding techniques, variational ansätze, and measurement in $\approx 500$ distinct model configurations. Our analysis reveals a clear hierarchy of influence on model performance. For hybrid architectures, which were benchmarked against their direct classical equivalents (e.g., the same architecture with the PQCs removed), the data encoding strategy is the dominant factor, with validation accuracy varying over $30\%$ for distinct embeddings. In contrast, the selection of variational ansätze and measurement basis had a comparatively marginal effect, with validation accuracy variations remaining below $5\%$. For purely quantum models, restricted to amplitude encoding, performance was most dependent on the measurement protocol and the data-to-amplitude mapping. The measurement strategy varied the validation accuracy by up to $30\%$ and the encoding mapping by around $8$ percentage points.
@@
I Introduction
In recent years, quantum machine learning (QML) has emerged as an interdisciplinary field at the intersection of quantum computing and machine learning Schuld2014; Wittek2014; Biamonte2017; Schuld2018. Its objective is to leverage phenomena in quantum mechanics to tackle problems from machine learning. A central component of many contemporary QML models is the parameterized quantum circuit (PQC), a flexible framework for creating trainable quantum circuits that can be used independently or integrated as hybrid components within classical neural networks (NNs).
This work investigates the use of PQCs for the task of satellite image classification using the EuroSAT dataset Helber2019. These PQCs serve as trainable quantum components that can be integrated into NNs or used by themselves. While PQCs offer a flexible paradigm for hybrid and purely QML models, their efficacy is highly sensitive to a multitude of design choices, with the most important being the encoding technique, the ansätze and the measurement strategy.
We study two different hybrid architecture networks and a purely quantum one, all inspired by convolutional neural networks (CNN). For the hybrid ones, we have a “Quantum-First” model, in which PQCs are used as convolutions to extract features, and a “Quantum-Later” model Senokosov2024, in which PQCs are used between the fully connected layers to process the extracted features Senokosov2024; Hur2022; Elhag2024. For the quantum architecture, we use three variations of PQCs as standalone components, one of which is inspired by the QCNN, introduced in Cong2019.
While the work of Bowles2024 provides a foundational benchmark focusing on the scaling of hyperparameters and the theoretical implications of entanglement, our work pivots to the structural composability of QML architectures. We argue that optimizing hyperparameters is complementary to selecting the correct combination of fundamental blocks. Therefore, rather than performing a hyperparameter grid search, we systematically evaluate the tripartite interaction between encoding, ansätze and measurement. Crucially, we analyze how this interplay varies across a spectrum of architectural paradigms, ranging from hybrid integrations (“Quantum-First” and “Quantum-Later”) to standalone QCNNs. This comprehensive scope allows us to distinguish between insights applicable generally to PQCs and those specific to structured quantum architectures, an angle not covered in Bowles2024.
@@ I Introduction
All our experiments were done through simulations using the CTIC Quantum Testbed (QUTE) platform qute. We used PennyLane bergholmPennyLaneAutomaticDifferentiation2022 to implement the PQCs, both in the purely quantum and hybrid models, and PyTorch paszke2019pytorchimperativestylehighperformance for the classical layers.
In the following Materials and Methods Section, we detail the dataset utilized for our experiments and explain the development process of the parameterized quantum circuits underlying our methodology. The Models Section presents the architectural designs investigated, along with the reasoning guiding their formulation. Empirical outcomes demonstrating model performance across diverse scenarios are summarized in the Results Section. Lastly, the Evaluation Section critically examines these findings, emphasizing key metrics and offering a balanced perspective on the efficacy and constraints of our approach.
II Materials and methods
II.1 Image classification
Our focus will be on classifying satellite images from the EuroSAT dataset Helber2019; examples can be seen in FIG. 1. EuroSAT was selected as our target due to its high intra-class variance and its standing as a standard, challenging dataset for evaluating remote sensing architectures. The dataset consists of $27,000$ geo-referenced images, provided by the Copernicus Earth observation program Copernicus using the Sentinel-2 satellite, and comes in two versions: RGB and all bands. The first consists of only RGB channels, and the other uses all the bands of the spectrum as collected from Sentinel-2. Each image channel has a resolution of $64\times 64$ pixels.
The classification task involves distinguishing between 10 distinct land-cover classes: annual crop, forest, herbaceous vegetation, highway, industrial, pasture, permanent crop, residential, river, and sea. The dataset is balanced, with $2000$ to $3000$ images per class, posing a significant challenge for models to identify specific textural and spatial patterns across diverse geographical regions.
In our experiments, we used the RGB version and downscaled them to $32\times 32$ and $16\times 16$. This downsampling represents a practical trade-off: satellite imagery at these scales preserves the essential spatial and spectral features required for land-cover classification. This is evidenced by the structural similarity index remaining above $0.9$ for the first case and over $0.78$ for the latter. Furthermore, this dimensionality reduction is vital for research in the current NISQ era; it ensures that the architectures remain computationally tractable for state-of-the-art simulators, allowing for the rigorous evaluation of hybrid models while the field transitions toward fault-tolerant hardware. Notably, the $32\times 32$ version was only studied for the hybrid models because of computation constraints. The $16\times 16$ version was studied for both hybrid and purely quantum models.
We selected this dataset as a continuation of the work conducted within the ARQA Cervera Technology Transfer Network ARQA. Nevertheless, identical experiments can be replicated on alternative datasets, provided they satisfy the previously established size constraints.
Figure 1: Samples from the EuroSAT dataset introduced in Helber2019. Each sample is a $64\times 64\times 3$ RGB image obtained from the Sentinel-2 satellite.
II.2 Parameterized quantum circuits
PQCs are the cornerstone of our approach. We integrate them as layers consisting of one or more circuits in the hybrid models and use them by themselves in the purely quantum ones. In general, PQCs consist of three key components: encoding, used to embed the information in a quantum state; variational ansatz, used to manipulate the encoded information; and measurement, used to extract information from the quantum state. FIG. II.2 shows the prototypical example of a PQC.
@@ II.2 Parameterized quantum circuits
Data encoding
As previously mentioned, the first stage in a PQC is the data encoding; this step serves as a mapping from data to a quantum state. Given that there is not a one-size-fits-all solution for the encoding step, we chose to include the following ones in our analysis:
  • Angle encoding Schuld2018; Schuld2014b; Cao2017: This is a prevalent and relatively straightforward method where $N$ classical data features, $x_{1},x_{2},\ldots,x_{N}$, are encoded into the rotation angles of $N$ individual qubits. Each feature $x_{j}$ is used as an angle parameter for a single-qubit rotation gate, such as $R_{X}(\theta_{j})$, $R_{Y}(\theta_{j})$, or $R_{Z}(\theta_{j})$, applied to a specific qubit. Notably, using $R_{Z}(\theta_{j})$ as an encoding gate acting on an initial state from the computational basis only adds an irrelevant global phase. The cost of this embedding is $\mathcal{O}(1)$.
  • Amplitude encoding Schuld2018: This technique embeds a classical data vector of $2^{n}$ normalized real numbers, $\{\alpha_{i}\}$, directly into the amplitudes of an $n$-qubit quantum state vector $|\Psi\rangle$. The state is represented as
@@
    where $|i\rangle$ are the computational basis states, and the amplitudes must satisfy the normalization condition $\sum_{i=0}^{2^{n}-1}|\alpha_{i}|^{2}=1$. The state preparation process required to achieve an arbitrary state $|\Psi\rangle$ can be computationally expensive, potentially demanding an exponentially deep circuit knill1995approximation.
@@
  • Instantaneous quantum polynomial-time (IQP) encoding Hav2019: This encoding extends the concept of angle encoding by incorporating entangling operations. $N$ classical features are typically encoded into $N$ qubits using single-qubit rotations, akin to angle encoding. These rotations are then interspersed with diagonal two-qubit gates, commonly $R_{ZZ}(\phi_{jk})$ gates, where $\phi_{jk}$ can be a function of individual features or products of features (e.g., $x_{j}x_{k}$). This structure allows for the generation of more complex quantum states that can potentially capture correlations between input features. The number of layers, $L$, is a tunable parameter. The scaling of this embedding is dominated by the two-qubit operations which in principle involve all pairs of qubits, thus making the scaling $\mathcal{O}(LN^{2})$.
@@
  • Quantum approximate optimization algorithm (QAOA) inspired encoding Lloyd2020: This encoding strategy draws inspiration from the structure of the QAOA. It transforms $N$ classical features into parameters that define a QAOA-like ansatz applied to $N$ qubits. The features typically determine the angles for rotation gates within layers corresponding to a problem Hamiltonian and a mixer Hamiltonian. The number of layers $L$ is a tunable parameter, and there are 3 variants: $X$, $Y$ and $Z$. The scaling of this embedding is dominated by the trainable two-qubit ZZ interactions, which connect nearest neighbors; hence the scaling is $\mathcal{O}(LN)$.
@@
  • Ring encoding Venturelli2023: This method combines angle encoding for $N$ features on $N$ qubits with a specific entangling topology. Following initial single-qubit rotations that embed the classical data, two-qubit entangling gates are applied between adjacent qubits arranged in a ring structure. Specifically, qubit $i$ is entangled with qubit $(i+1)\mod N$; thus, $N$ two-qubit entangling gates are applied. The number of layers, $L$, is a tunable parameter. Since consecutive entangling gates share one qubit, the cost of ring encoding is $\mathcal{O}(NL)$.
@@
  • Waterfall encoding Venturelli2023: Similar to Ring encoding, Waterfall encoding also utilizes angle encoding for $N$ features on $N$ qubits, followed by a structured application of entangling gates. In this scheme, the entangling gates are typically arranged in a “waterfall” pattern. For example, qubit $i$ might control an operation on qubit $i+1$, which in turn might control an operation on qubit $i+2$, and so forth. In each layer, there are N(N1)/2N(N-1)/2 entangling gates; hence, worst case scaling is $\mathcal{O}(LN^{2})$.
Variational ansatz
@@
The variational circuit constitutes the second part of a PQC. It consists of a set of quantum gates with trainable parameters, akin to neurons in a NN. Its purpose is to manipulate the quantum state generated by the previous step with the goal of turning it into a state whose measurement, in a given basis, solves the task at hand. We chose to experiment with the following ones:
@@
  • No entanglement: This circuit uses randomness in the assignment of quantum gates, applying either an RZ gate or the identity operation to each qubit in a stochastic manner. Due to the potential presence of trivial operations (identity), the resulting circuit typically exhibits a lower computational depth compared to more densely parameterized or entangled circuits.
@@
  • Full entanglement Schuld2020: This circuit is implemented using the StronglyEntanglingLayers module from PennyLane bergholmPennyLaneAutomaticDifferentiation2022. It allows for a configurable number of layers, each consisting of a sequence of single-qubit rotations followed by entangling operations. This structure enables flexible depth and entanglement patterns suitable for various variational quantum algorithms.
    Figure 3: Visualization of Full entanglement ansatz.
@@
  • Ring circuit Venturelli2023: It performs single-qubit rotations using RY and RZ gates, followed by the application of controlled-Z (CZ) gates arranged in a ring topology. The resulting structure resembles the ring embedding, but with entanglement introduced through CZ operations instead of CNOT.
    Figure 4: Visualization of Ring ansatz.
@@
  • NQ circuit Hur2024: The circuit consists of multiple layers, each combining parameterized single-qubit rotations, RY and RZ, followed by entangling gates (CNOT) in a linear nearest-neighbor topology.
    Figure 5: Visualization of NQ ansatz.
@@
  • QCNN circuit Cong2019: circuit structure inspired by classical CNNs in which convolutional and pooling operations are subsequently applied in each layer (see FIG. 7(c)). In our implementation, when the number of qubits at a layer is odd, one is excluded from the pooling, although it is still considered in the convolutional stages. At the end of the circuit, the qubits that were discarded are recovered for the fully connected layer. Given that the number of free parameters in an arbitrary unitary scales as $4^{N}-1$, we substituted this last operation by the Simplified Two Design cerezo2021cost.
Measurement
@@
  • Z-basis: It is also called the computational basis. This is the most commonly used measurement in quantum circuits because it collapses a qubit into $|0\rangle$ or $|1\rangle$.
@@
  • Histogram: We measure each qubit on the Z basis and count the frequency of each of the possible $2^{N}$ outcomes, for $N$ qubits. Each state is associated with a class; hence, the model in this case will try to learn to link classes with corresponding states. This is in essence a Quantum Circuit Born Machine liu2018differentiable in which we set the most frequent sample to be the predicted class label.
  • Pauli string (Paulis): We estimate the expectation value of Pauli strings of $N$ qubits, i.e., observables of the form $O=\otimes_{i=0}^{N}P_{i}$, where $P_{i}$ is a Pauli gate acting on the ithi-th qubit. Here, a class is identified with an operator such that the model is trained to maximize the expectation value of that observable when images from the correct class are fed into the model.
Recent studies have also pointed out the relevance of the type of measurement considered in the models chen2025learningprogramquantummeasurements.
III Models
For an overview of the models, see TABLE 1. This table outlines the distinct components integrated into each model, including data encodings, ansätze, and measurement strategies.
III.1 Hybrid models I: HQNN-Parallel
The HQNN-Parallel was first introduced in Senokosov2024 (shown in FIG. 7(a)). This architecture introduces a NN layer based on multiple PQCs. From this point onward, we will call this a quantum linear layer due to its similarity with a classical linear layer.
A quantum linear layer operates similarly to its classical counterpart. It takes in an input vector of size $V$ and outputs another of size $W$. However, unlike the classical case, here $W\leq V$ given that at most we do a single measurement on all qubits. In the classical version, the linear layer is defined by the number of input and output features. The quantum analog is characterized by the number of input features, the number of qubits allocated per circuit (the same number for all), and the encoding used, which will determine the proportion between input and output features.
In the simplest case, we have an input vector of length $L$, we allocate $L/2$ qubits per circuit, and we use an encoding that maps a single feature to a single qubit, like angle encoding. In this case, the quantum linear layer will consist of two PQCs of $L/2$ qubits each, and the output will be a vector of length $L$, given that we measure all qubits on a given basis. This scheme is shown in FIG. 6. For instance, if we use amplitude embedding, then each $N$ qubit circuit will process $2^{N}$ features, and the output will be of size $N\times\text{the number of circuits}$.